OpenStudy (anonymous):

last year an automotive technician has a gross income of 32,200 which she contributed 6% to her 401(k)plan if she got paid weekly how much was deducted for each pay check for her 401(k)plan

OpenStudy (anonymous):

I got 27600

OpenStudy (anonymous):

\[\frac{32200}{52}*\frac{6}{100}\text{=}\frac{483}{13}=\text{$\$$37}.15 \text{ per} \text{ week}\]

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

You're welcome.